OpenStudy (anonymous):

Janine made a cylindrical vase in which the sum of the lateral area and area of one base was about 3000 square centimeters. The vase had a height of 50 centimeters. Find the radius of the vase. Explain the method you would use to find the radius.

OpenStudy (anonymous):

height of the cylindrical portion h=50-r (r radius of base) as per the prob 2(pi)(r)(50-r)+(pi)(r)(r)=3000 or (pi)r(100-r)=3000 this being a quadratic can be solved

OpenStudy (anonymous):

how long u gonna be on here because i have to take my son to school

OpenStudy (anonymous):

Its 5:34am ill be on till 12:30 here so 7 hours.